In a sale, the price of a computer is reduced from $1,280 to $1,030. Find the rate of the discount and the amount of the discount. Round the discount rate to the nearest tenth of a percent if necessary.
Answer by TimothyLamb(4379)

discount = 1280 - 1030 = $250
---
rate = ( 250/1280 ) * 100 = 19.5%
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php